OpenStudy (anonymous):

Enter the missing exponent (17^4)^5 = 17 what is the value of 5*6 - 2[8 +(12 - 6)] simplify (4^5 * 4^-3)^2/4^2 @RyleeC these are the last quistions

OpenStudy (anonymous):

For 1 i think the missing ^ number is 9

OpenStudy (anonymous):

1. 20 because a power to power is times. you add them when you multiply variables with different powers.

OpenStudy (anonymous):

2. you can do that just solve the insides first than multiply

OpenStudy (anonymous):

3. idk because I forgot how to do negative powers

OpenStudy (anonymous):

the answer for 2 is ":two"?? because 12-6 = 6 and 8+6 = 14and 14 * 2 = 28 Then 5*6= 30 and 30-28=Two"2"???

OpenStudy (anonymous):

@RyleeC ^^

OpenStudy (anonymous):

its 5^6-28 I cant do 5^6 without a calculator.